OpenStudy (anonymous):

A line passes through (2,-1) and (8,4) a. write an equation for the line in point-slope form b. rewrite the equation in standard form using integers Please help me find the answer?

OpenStudy (anonymous):

Well, first find the slope, so you need to do with to points (a,b) and (c,d) To get: m = (d-b)/(c-a) Then, you have m, so you put in the point slope equation like so: y = mx +b. Then use one of your points for say x and y, like the point (8,4) And put in for those variables: So you get: 4 = m*8+b And solve for b, now that you have b and m, you can put your final equation in the form of: y=mx+b (b and m should be numbers we found, leave y and x as a variables this time. So what do you get for your equation? Do this, then we'll do part b.

OpenStudy (anonymous):

y-1=5/6(x-2) ??

OpenStudy (anonymous):

No, remember we want point slope form, so do like I showed you in the steps first.